OpenStudy (anonymous):

Solve each system using elimination. tell wether the system has one solution, infinitely many solutions, or no solution. 5x+4y=-83 3x-3y=-12

hero (hero):

y = (-83-5x)/4 y = (12 + 3x)/3 y = y 4(12 + 3x) = 3(-83 - 5x) 48 + 12x = -249 - 15x 27x = -297 x = -297/27 x = -11 y = -7

hero (hero):

The system has only one solution